[figure: 9.7 cm]

Specific heat some of the R_{3}Co compounds.

N.V.Tristan^{a}, S.A.Nikitin^{a,b}, T.Palewski^{a}, K.Nenkov^{a,c}, K.P.Skokov^{d} aInternational Laboratory of High Magnetic Fields and Low Temperatures, 95 Gajowicka str., 53-421 Wroclaw, Poland

bFaculty of Physics, Moscow State University, Vorobievy Gory, 119899 Moscow, Russia

cInstitut fur Festkorper-und Werkstofforschung Dresden, Postfach 270016, 01171

Dresden, Germany

dFaculty of Physics, Tver State University, 33 Zheljabova str., 170002 Tver, Russia

Specific heat measurements were performed in the temperature range from 2 to 300 K on the number of the R_{3}Co compounds. Using the value of the electronic specific heat coefficient γ=0.017 J/(mol K) and Debye temperature Θ_{D}=234 K for Y_{3}Co compound, the Debye temperatures, phonon and conduction electronic contributions as well as magnetic part of the specific heat were estimated for other R_{3}Co compounds. The magnetic part of the specific heat can be divided into two parts, one of which caused by entropy change due to magnetic ordering lower Neel temperature T_{N} and the other part is due to Schottky anomaly with broad maximum around 26-36 K depend on rare earth component of the compounds. Magnetic contribution to the entropy S_{m} was calculated using experimental data. The value of the magnetic entropy S_{m} does not reach theoretical one S_{m}^{max} at T=T_{N}. For Gd_{3}Co the expected value of the entropy is reached at about room temperature while behavior of the S_{m}(T) for Er_{3}Co and Ho_{3}Co compounds is different from expected. The magnetic contributions to entropy caused by localized 4f and itinerant 3d electrons are discussed.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : N.V.Tristan

Address for correspondence : International Laboratory of High Magnetic Fields and Low Temperatures 95 Gajowicka str. 53-421 Wroclaw Poland

Email address : lnatalka@alpha.mlspmint.pan.wroc.pl